OpenStudy (anonymous):

Which expression is equivalent to this one? u • 5 + u • v A. u(5 + v) B. 5(u + v) C. v(5 + u) D. u · (5 + u) · v

OpenStudy (anonymous):

@dtan5457 @DanteFemboy

OpenStudy (anonymous):

Which one do you think it is?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

5(u + v) = 5 * u + 5* v does that equal u • 5 + u • v ?? You were close though

OpenStudy (anonymous):

Want to give it another try?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Correct!!!

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

Sorry it is x = 8 I made a mistake. Ok, we need to solve for x 5(8 + 7) = 5 • x + 5 • 7 First do what is inside of the parenthesizes(8+7) 5(8 + 7) = 5 • x + 5 • 7 5(15) = 5 * x + 5 * 7 Next times everything together 5(15) = 5 * x + 5 * 7 75 = 5x + 35 Now lets get 5x by its self by using the addition property. 75 = 5x + 35 75 - 35 = 5x 40 = 5x Next use the division perperty to get x alone. Divide by 5 on both sides of the = 40 /5 = 5x/5 8 = x x= 8

OpenStudy (anonymous):

thx